OpenStudy (anonymous):

A square photograph with sides of 15 inchs must be ruduced o an area 40% of its orginal size to it in a book. Find the length of each side of the reduced photograph. Round to nearesst of the tenth

OpenStudy (anonymous):

15^2 = A A = 225 inch.^2 .40 * 225 = 90. \[\sqrt{90} = 9.4868\] So 9.5 is the side length

OpenStudy (anonymous):

why wouldnt you do 225-90 and then get the sqare root of that

OpenStudy (anonymous):

Because you said you want an area of 40% it's original. If u had wanted to only reduce 40% then you would do 225-90.